rounded to the nearest tenth, algebraically find the solution(s) of logx+log2x=1
Answer by nerdybill(7384)

rounded to the nearest tenth, algebraically find the solution(s) of logx+log2x=1
.
log(x * 2x) = 1
log(2x^2) = 1
2x^2 = 10^1
2x^2 = 10
x^2 = 5
x = sqrt(5)
x = 2.2 and -2.2